OpenStudy (karatechopper):

One number is 5 times another number. The product of the two numbers is 245. Find the two numbers. Please explain this to me!!!

OpenStudy (anonymous):

You're right about the equations 5N = X and NX = 245. What I would do is substitute 5N for X in the second equation since you know one X is worth 5N. So you would get: N(5N) = 245; then distribute to get: 5N ^2 = 245; then divide both sides by 5 to isolate the variable N^2 = 49; then take the square root of both sides N = 7; put that value back into your original equation to get X = 5N = 5(7) = 35 N could also be -7 since both 7 and -7 square to be 49. In this case, X would be -35.

OpenStudy (anonymous):

If 5n = x, then put 5n in for x in the other equation 5n^2 = 245 then divide by 5 n^2 = 49 n = 7 5 times n = 35 The 2 numbers are 7 and 35

OpenStudy (karatechopper):

see i did the substitute part then got confused

OpenStudy (anonymous):

At least you tried